OpenStudy (anonymous):

1 triangle side length is 2 and the other one is 2. What is the 3rd side length? My choices are 3,4, and 5.

OpenStudy (anonymous):

It's either a equilateral or right triangle.

ganeshie8 (ganeshie8):

you may use this :- \(\large \text{third side must always be less than sum of the other two sides}\)

ganeshie8 (ganeshie8):

since you're given two sides of lengths :- 2, 2 add them, : 2+2 = 4

ganeshie8 (ganeshie8):

So, third side must be less than 4. look at ur choices, which one is less than 4 ?

OpenStudy (anonymous):

I'm kinda confused. I thought we were supposed to square them?

ganeshie8 (ganeshie8):

see my first reply

ganeshie8 (ganeshie8):

We cannot use squaring here cuz, squaring is allowed oly in Right triangles.

ganeshie8 (ganeshie8):

my first reply is also called 'triangle inequality theorem'

OpenStudy (anonymous):

Lololol still confused. I did this : 2x2=4 +2x2=4 =8. My choices I may pick from are 3,4, and 5. And I squared each of those and they are all greater than 8.

OpenStudy (anonymous):

OHHHHHHH I GET IT NVM

ganeshie8 (ganeshie8):

we dont need to all that circus :) did u take time to read my replies above ? :)

ganeshie8 (ganeshie8):

we just need to add the side lengths 2+2 = 4 Since, 3rd side must always be less than this sum, 3rd side < 4 u have oly one option satisfying this inequality !

ganeshie8 (ganeshie8):

glad to hear u get it ! :D